remi2019:

A hyena can run one over four of a mile in 22.5 seconds. At this rate, which equation can be used to determine how fast a hyena runs in miles per hour?

remi2019:

plzzz help

supie:

Choices?

remi2019:

? one-fourth mile over twenty-two point five seconds = 40 miles per hour twenty-two point five seconds over one-fourth mile = 40 miles per hour one-fourth mile over twenty-two point five seconds = 90 miles per hour twenty-two point five seconds over one-fourth mile = 90 miles per hour

remi2019:

1/4 mile/22.5 sec = 40 22.5 sec/1/4 mile = 40 1/4 mile/ 22.5 sec = 90 22.5 sec/1/4 mile = 90